OpenStudy (anonymous):

whats the interest on $175.00 at 4% for 120 days

OpenStudy (anonymous):

P=175 R=4% T=120 days Interest=P*R*T Interest=175*(4/100)*120/365 Interest=$2.3013

OpenStudy (jdoe0001):

\(\bf i = prt \\ \quad \\ i = interest\\ p=principal \to 175.00\\ r = rate\to 4\%\to \frac{4}{100}\to 0.04\\ t=years\to 1year=365days\quad thus\quad \frac{120}{365} \approx 0.3287year \\ \quad \\ i = prt\implies 175.00\cdot 0.04\cdot 0.3287\)